OpenStudy (anonymous):

describe the form of the limit. Does L'Hopital's rule apply? a. 0/0 b. infinity/infinty c. infinity - inifinity d. infinity*0 e. 1 ^ infinity f. zero^ zero g. infinity ^ zero h. the form limit is none of these please remember to say yes or no if as to whether the rule applys ..thank you!

OpenStudy (anonymous):

my educated guess. yes to only: a, b, e

OpenStudy (anonymous):

a and b

OpenStudy (anonymous):

L'Hopital's Rule only applies to a form of 0/0 or infinity/infinity. Howeber, other forms can be changed to 0/0 or infinity/infinity. But the Rule can only be applied when it is of the form 0/0 or infinity/infinity. Some of your other forms are indeterminate forms, and can be changed to 0/0 or infinity/infinity. Some cannot, and may involve taking the ln of the expression, etc.

OpenStudy (anonymous):

i think 1^inf as well

OpenStudy (anonymous):

You cannot use direct L'Hopital's Rule with 1^infinity.

OpenStudy (anonymous):

but you would have interest in turning it into a 0/0 or inf/inf form because it is identified as 1^inf like: \[\lim_{x-> \infty} \left( e^{\frac{ 1 }{ x }} \right)^x\]

OpenStudy (ranga):

For 1^infinity, you have to take ln first. For example, lim x->0 (1 + x)^(1/x) Can't apply L'Hopital's Rule directly. y = (1 + x)^(1/x) Take ln on both sides and then apply L'Hopital's Rule.

OpenStudy (anonymous):

Example; the function (sec x)^ (cot x) as x---->0 is of the form 1^infinity...but you cannot use a direct L'Hopital's Rule. As I said earlier, you would try to convert it, if possible, to 0/0 or inifinity/infinity.

OpenStudy (anonymous):

but you knew to that because it was 1^inf

OpenStudy (anonymous):

So I agree with @ranga

OpenStudy (anonymous):

my example is silly. e^(1/x)^x = e . lol forget it

OpenStudy (anonymous):

agree @Euler271

OpenStudy (anonymous):

i agree

OpenStudy (kainui):

Classic example, \[\lim_{x \rightarrow \infty} (1+\frac{ 1 }{ x})^x\] appears to be 1^infinity, but is equal to e.